OpenStudy (anonymous):

how to find x-intercept and y-intercept? example: 2 x squared + 3 y squared = 18 .

OpenStudy (anonymous):

first put x equals 0

OpenStudy (anonymous):

and substiute in th equation gives you

OpenStudy (anonymous):

3y^2 equals 18

OpenStudy (anonymous):

x-intercept means the point where curve crosses x axis so put y=0 in the given equation and find x intersept and for y intercept put x=0 to find the point where curve crosses y axis

OpenStudy (anonymous):

y^2 equals 6

OpenStudy (anonymous):

ahh.. got it. :) thank you. ^_^

OpenStudy (anonymous):

ok gd

OpenStudy (anonymous):

this for y intercept

OpenStudy (anonymous):

and y equals 0 for the x intercept

OpenStudy (anonymous):

eh how about, y = x^2 - 3x + 2, what is the y-intercept?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

since x equals 0 and subs.

OpenStudy (anonymous):

ahh... got it again.. :) eh how about, y = x^3 - 2x^2 + x - 2 , what is the x-intercept?

OpenStudy (anonymous):

x equals 2

OpenStudy (anonymous):

take y equals 0 first

OpenStudy (anonymous):

x^3 - 2x^2 +x-2 equals zero then take common factor x^2

OpenStudy (anonymous):

x^2(x-2)+x-2 equals 0

OpenStudy (anonymous):

x^2 (x-2) + 1 (x-2) = 0 .

OpenStudy (anonymous):

take common x-2

OpenStudy (anonymous):

x^2 + 1 (x-2) = 0.

OpenStudy (anonymous):

no (x-2)(x^2+1)

OpenStudy (anonymous):

ahh.. got it again. :) hihi.. thanks a lot. :) eh, how about, how to know the symmetry of a function?